solve x+1/3+x+6/7=1
Answer by checkley71(8403)

X+1/3+X+6/7=1 OR 2X=1-1/3-6/7 OR 2X=1-(7+18)/21 OR 2X=1-(25/21) OR 2X=-4/21 OR
X=(-4/21)/2 OR X=-4/42
PROOF -4/42+1/3-4/42+6/7=1 OR -4/42+14/42-4/42+36/42=1 OR -8/42+50/42=1 OR
42/42=1 OR 1=1